OpenStudy (anonymous):

A person is standing 300 feet away from a building. The angle of elevation from the person to the roof of the building is 49 degrees. Find the height of the building to the nearest foot.

OpenStudy (lgbasallote):

ahhh..i'll let @coco answer this one ^_^

OpenStudy (anonymous):

1. - With a scale of 1cm : 3ft .... ==> 17 feet is represented by 17/3 cm on the drawing .... 17/3 = 5 2/3 = 5.7cm (to nearest tenth) ==> 10 feet is represented by 10/3 cm on the drawing .... 10/3 = 3 1/3 = 3.3 (to nearest tenth) 2. - By using what we know about the geometry of a rectangle, with b base length, and h height length, and p defined as perimeter .... ==> 2b + 2h = p .... but they tell you p = 48 so we have .... ==> 2b + 2h = 48 Then just feed in each of the 'b' values supplied and re-arrange to find the value of h .... ==> 2(6) + 2h = 48 ==> 2h = 48 - 12 ==> h = 36/2 = 18 units. ==> 2(8) + 2h = 48 ==> 2h = 48 - 16 ==> h = 32/2 = 16 units. ==> 2(12) + 2h = 48 ==> 2h = 48 - 24 ==> h = 24/2 = 12 units. ==> 2(16) + 2h = 48 ==> 2h = 48 - 32 ==> h = 16/2 = 8 units. 3. - This question defines a RAT (Right Angled Triangle - or Right Triangle for our USA buddies) with angle 49, (A)djacent side 300ft, and they want the (O)pposite side. Now, in the UK students learning this stuff are offered the following Mnemonic to aid choice of trig equation .... SOH - CAH - TOA - you just tick off the 2 things you know about the triangle and the one thing you are seeking, i.e. here we know S & C & T (in other words we can calculate Sin49, Cos49, and Tan49 if needed), we also know A = 300 feet, and we want O .... so we use the Tan equation :o) ==> Tan49 = O/300 ==> O = 300Tan49 ==> O = 345.1105222 = 345 feet (to nearest foot)

OpenStudy (anonymous):

Let h be the height of the building  h/300 = tan 49 (h/300)(300) = (tan 49)(300) h=345 feet